# Towards Browser Controls to Protect Cookies from Malicious Extensions

### v2

@@ 3.1 Manipulating Cookies via Extension APIs
    Case 1 - Cookies API Extension: This extension implements a cookie management interface allowing the user to view, edit, and delete browser cookies. When fetching cookies the extension also forwards them to a malicious server. Using this we stole the cookies of many websites including Google, Amazon, and Facebook. Adding the stolen cookies to a new Firefox instance allowed us to access the corresponding test accounts without their credentials. We note that these accesses require a combination of cookies rather than a single cookie, but as extensions can easily access all of a website’s cookies this is not an obstacle.
@@
    Case 2 - WebRequest API Extension: This extension allows the user to intercept, view, and modify traffic for specified websites. When viewing the request, the request headers are also forwarded to a malicious server. From the headers, we can steal all the cookies sent to and set by the website. This extension can also modify request headers and thus edit the cookies sent to/received from the server. It is important to reiterate that modifying headers requires the WebRequestBlocking permission [59] which is only available for V2 extensions in Chrome. While V3 extensions cannot edit requests they can still steal cookies.
@@
    Case 3 - DeclarativeNetRequest API Extension: This extension allows the user to create rules that modify network traffic without viewing its content. As such, it allows us to edit the cookies in traffic, however, as the extension cannot inspect the request headers, it cannot steal cookies. The Declarative API only exists for V3 extensions.
Table 1: Breakdown of extensions versioning, host permissions, and access to cookies
@@
Table 2: Total users affected by each API
@@
Table 3: Extensions with "$<$all_urls$>$" and cookie access
@@
Most modern browsers are implemented on top of a few open-source browsers. Chrome, Microsoft Edge, Opera, and Brave are all built on top of the Chromium browser [38]. As such, our extensions worked on all four browsers without modification. We also tested the extensions in the Firefox-based browsers: Firefox and Waterfox. The V2 extensions (Cookies and WebRequest) were directly transferable as Firefox also implements the WebExtension APIs [62] to support cross-browser extensions. When using V3, one change was required. Despite requiring that host permissions are defined in an extension’s manifest, Firefox treats them as optional permissions for V3 extensions [63]. Thus, after installation, the extension must prompt the user to re-accept its host permissions. Once granted, V3 extensions behaved as they did in Chromium. We believe that prompting the user does not pose a problem to the adversary, as they could (and likely would) imitate a legitimate service. Another caveat is that Firefox still supports the WebRequestBlocking permission for V3 extensions. Testing in Waterfox obtained the same results.
3.2 Extension API Permissions’ Landscape
@@
The above examples highlight how certain extension APIs can be leveraged to steal and modify cookies (including session cookies). To determine the prevalence of these permissions we crawled Chrome’s and Firefox’s extension repositories: the Chrome Web Store [64] and addons.mozilla.org (AMO) [65]. As of March 2024, we identified 122,951 Chrome and 37,783 Firefox extensions. However, of the extensions identified, we could only pull the manifests of 119,152 Chrome and 37,631 Firefox extensions for analysis. Appendix A discusses the scrape and identified extensions in greater detail. We also identified several Chrome Apps and Android extensions which we discuss in Appendix B. All extensions discovered and manifests analyzed are available in our public repository [39].
@@
We parsed each extension’s manifest for three attributes. First, we obtained the manifest version from the manifest_version tag. Next, we checked the permissions entry to find if the extension used any of the three APIs discussed in Section 3.1. Finally, we searched the manifest for the strings “$<$all_urls$>$”, “*://*/*”, “https://*/*”, and “http://*/*”. These strings allow an extension to access every page (first two strings), all HTTPS pages, and all HTTP pages in the browser, respectively [66]. As these host permissions can appear in various places, their presence indicates that at least a portion of the extension has this access. Similarly, not all extensions with “$<$all_urls$>$” behavior use this tag. Due to this, we treat “$<$all_urls$>$”, “*://*/*”, and the combination of “https://*/*” and “http://*/*” as equivalent. Table 1 summarizes our results. We also calculate the number of users affected by each API as reported by each extension store (depicted in Table 2).
@@
We found that 59.9% of extensions are written using manifest V2. However, V2 extensions are much more prevalent in Firefox accounting for 92.8% of all Firefox extensions. Chrome extensions are more evenly split with 49.5% still using the V2 format. This disparity is likely due to Chrome’s plan to deprecate the V2 format in 2025 [67] whereas Firefox will continue to support both versions [68].
@@
For host permissions, we found that 35.2% of the extensions analyzed define broad host permission. Unlike manifest versioning, this ratio is consistent across Firefox (35.6%) and Chrome (35.3%). Nearly all of these extensions have the “$<$all_urls$>$” behavior. This accounts for 34.2%, 35%, and 33.9% of all, Firefox, and Chrome extensions, respectively. 0.2% of extensions declared permission for all HTTP sites only and 1% declared permission for all HTTPS sites alone.
@@
7,882 extensions (5% of extensions analyzed) used the Cookies API, reaching a combined user base of 422,847,233 users. Note that this combined user base represents an upper bound, as a single user may have multiple extensions and thus be counted multiple times. Similarly, it is unclear whether extension repositories measure users by installation or actual user accounts. More extensions used the WebRequest API. Specifically, 11,839 (7.6%) were able to monitor network traffic and 6,978 extensions (4.5%) also declared the WebRequestBlocking permission and could edit HTTP headers. Again while limited extensions declare these permissions, they have nearly one billion users (994,703,506) with 533,244,862 users simultaneously affected by the WebRequestBlocking permission. The DeclarativeNetRequest API is the least used, only appearing in 2,114 extensions (1.3%). Similarly, it is almost non-existent in Firefox appearing in only 66 extensions (0.17% of all Firefox extensions). This low adoption is likely due to the ongoing transition from V2 to V3 formatting (at least in Chrome), as the DeclarativeNetRequest API is only available to V3 extensions. In that regard, it appeared in 3.4% of V3 extensions found and has a combined 353,734,697 users.
@@
Each API is limited by the extension’s host permissions. Hence, we also examined how many of the above extensions also have “$<$all_urls$>$” host permissions. Table 3 summarizes these findings. 54.8% of extensions with Cookies API permissions have access to all cookies in the browser. Similarly, 65.2% of WebRequest extensions and 63.7% of WebRequestBlocking extensions can read and modify any requests in the browser. 67.9% of manifests with DeclarativeNetRequest permissions can also modify any network requests. We also found that 21.3% of these extensions (14.5% all declarative extensions) also contain WebRequest permissions, allowing them to view any request (despite contradicting the privacy-preserving philosophy of the declarative API).
@@
From our analysis, we find that widespread extensions with a combined user base in the order of hundreds of millions can access every cookie in the user’s browser. When compromised (or malicious), these extensions pose a significant threat to cookies and session security.
4 CREAM Overview
@@
Motivated by the results presented in Section 3, we advocate for browser controls to allow selective prevention or detection (when prevention is not possible due to application requirements) of cookie theft and illegal cookie modifications made by extensions. To realize these goals we propose CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modification to the trust model within browsers. This Section focuses on CREAM’s concepts and protocol. We discuss browser support for CREAM in Section 5.
4.1 Threat Model
We extend the active network attacker model described in prior work [15, 16, 19]. The model assumes an Adversary ($\sf{\mathcal{A}dv}$) can intercept, modify, and inject traffic on the network. It also assumes $\sf{\mathcal{A}dv}$ controls a malicious website. In addition, we assume that $\sf{\mathcal{A}dv}$ controls a malicious extension in the user’s browser with arbitrary API permissions and can run on all pages in the browser (e.g. has the “$<$all_urls$>$" host permission). As such, $\sf{\mathcal{A}dv}$ can also inject content scripts (malicious JS) into all pages. $\sf{\mathcal{A}dv}$’s goal is to modify or steal a user’s cookies through attacks such as cookie hijacking and cookie fixation and gain access to the user’s session.
4.2 Rationale
CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly is a 1-byte flag indicating whether a cookie should be visible to extensions. If set, the cookie is only accessible to the browser. Therefore neither JS nor extensions can access these cookies. The BrowserOnly attribute only protects cookies within the browser. To prevent BrowserOnly cookies from being stolen on the network they should also be marked Secure, which is standard.
For cases where cookies must be accessible by websites and extensions (where BrowserOnly can not be used), we define the Monitored attribute. The Monitored attribute holds a server-defined message that when received is stored by and only accessible to the browser. The browser also logs all changes made to Monitored cookies. Later, when sending a Monitored cookie to the server, the browser attaches an authenticated report to the request that contains this server-defined message, the cookie’s current settings, and any logged changes. With this, the server can validate the changes made to the cookie and detect if it is stolen.
Since Monitored cookies must be accessible, we cannot prevent them from being stolen. Instead, we make stolen cookies unusable (e.g., for authentication). Figure 2 depicts the protocol to create (part (A)) and use (part (B)) Monitored cookies. The browser generates a symmetric key for each site it visits (step A.1 in Figure 2) and attaches it to all outbound HTTPS requests for that site (step A.2). Note that in this context, a site refers to the protocol and registry controlled domain name of a resource [17, 69] (e.g., https://example.com). The generated key allows a site to identify the browser in subsequent communication. However, by creating a new key per site these keys cannot be used to track users across different sites. When creating a Monitored cookie, the server builds the Monitored attribute’s message. This message must include the cookie’s name, value, and the received browser key. The cookie’s name-value pair ties the attribute to this specific cookie while the browser key ties the whole cookie to a single browser. Along with these values, the server can include any other data necessary for later cookie validation. The server then uses a secret to symmetrically encrypt the message and uses the resulting ciphertext as the Monitored attribute’s value (step A.3). This authenticated encryption ensures the attribute was generated by the server, is only readable to the server, and cannot be altered.
@@ H $=HMAC(Report_{C_{i}},K_{browser})$
When sending the Monitored cookie back to the server, the browser builds its corresponding Monitored report, including an HMAC produced with the generated key for that site, and attaches it to the request (steps B.1-B.3). Upon receipt, the server decrypts the Monitored message with its secret key (B.4), validates that it is the correct message for the received cookie (B.5), and verifies the HMAC using the key stored in the Monitored attribute (B.6). If the cookie was stolen by a malicious extension, the HMAC verification would fail as $\sf{\mathcal{A}dv}$ would not know the browser’s key (as long as the browser is itself trusted – recall Section 4.1). Similarly, attempts to change the server-defined Monitored message would cause the decryption/validation to fail. As a result, Monitored cookies can only be used by the legitimate browser.
As detection relies on the symmetric key sent to the server in the initial request, the key can only be sent over secure (HTTPS) channels. This ensures the key is always encrypted and inaccessible to $\sf{\mathcal{A}dv}$ over the network. Otherwise, if the key was not encrypted $\sf{\mathcal{A}dv}$ could steal it to forge authenticated Monitored reports and use stolen Monitored cookies. However, once established, Monitored cookies can be sent over secure and insecure channels as the key in the Monitored attribute is always encrypted by the server’s key.
5 CREAM Architectural Design
As discussed above, the BrowserOnly and Monitored attributes require architectural support from the browser to restrict/monitor extension actions. Aside from creating the new cookie attributes, this also requires support from the Cookie Monster and Network Service. As each attribute functions independently, we discuss their required support separately.
@@ 5 CREAM Architectural Design
5.1 BrowserOnly Attribute
As discussed in Section 4, BrowserOnly is a 1-byte flag that prevents a cookie from being accessed by JS and extensions. Thus the Cookie Monster must be updated to filter on this attribute. Recall from Section 2.3.2 that there are two ways to fetch cookies from the Cookie Monster, GetAll and GetList, and both GetList and Set filter cookies based on a supplied set of options. As such, we create a new option to exclude BrowserOnly cookies from GetList and Set calls. This option prevents BrowserOnly cookies from being returned or overwritten in most cases. However, as GetAll never filters its results, BrowserOnly cookies are still returned. To address this, we create a new endpoint (GetAllExt) that returns all non-BrowserOnly cookies in the Cookie Monster. These changes are then propagated to the Cookie Manager API and its respective implementations. Finally, the Cookies extension API is updated to use GetAllExt instead of GetAll and the option to exclude BrowserOnly is added to the necessary accesses. This blocks all extension and JS accesses to BrowserOnly cookies through the Cookie Monster. Note that cookie deletion in response to storage restrictions (“purges”) can be conditioned on cookies’ priorities and security attributes [58]. Therefore, we also configure the Cookie Monster to not evict BrowserOnly cookies in these scenarios, thus mitigating eviction attacks [5, 70].

### v3

@@ III-A Manipulating Cookies via Extension APIs
    Case 1 - Cookies API: This extension implements a cookie management interface allowing the user to view, edit, and delete browser cookies. When fetching cookies, the extension also forwards them to our test server. Using this, we captured the cookies for many websites including Google, Amazon, and Facebook. Adding the stolen cookies to a new Firefox instance allowed us to access the corresponding test accounts without their credentials. We note that these accesses require a combination of cookies rather than a single cookie, but as extensions can easily access all of a website’s cookies this is not an obstacle.
@@
    Case 2 - WebRequest API: This extension allows the user to intercept, view, and modify traffic for specified websites. When viewing the request, the request headers are also forwarded to our test server. From the headers, we can retrieve all the cookies sent to and set by the website. This extension can also modify request headers and thus edit the cookies sent to/received from the server. It is important to reiterate that modifying headers requires the WebRequestBlocking permission [65] which is only available for V2 extensions in Chrome. While V3 extensions cannot edit requests they can still steal cookies.
@@
    Case 3 - DeclarativeNetRequest API: This extension allows the user to create rules that modify network traffic without viewing its content. As such, it allows us to edit the cookies in traffic, however, as the extension cannot inspect the request headers, it cannot steal cookies. The Declarative API only exists for V3 extensions.
TABLE I: Breakdown of extensions versioning, host permissions, and access to cookies
@@
TABLE II: Total users affected by each API
@@
TABLE III: Extensions with ”$<$all_urls$>$” and cookie access
@@
Most modern browsers are implemented on top of a few open-source browsers. Chrome, Microsoft Edge, Opera, and Brave are all built on top of Chromium [42]. As such, our extensions worked on all four browsers without modification. We also tested the extensions in the Firefox-based browsers: Firefox and Waterfox. The V2 extensions (Cookies and WebRequest) were directly transferable as Firefox also implements the WebExtension APIs [68] to support cross-browser extensions. When using V3, one change was required. Despite defining the host permissions in each extension’s manifest, Firefox treats them as optional permissions for V3 extensions [69]. Thus, after installation, the extension must prompt the user to re-accept its host permissions. Once granted, V3 extensions behaved as they did in Chromium. We believe that prompting the user does not pose a problem to the adversary, as they could (and likely would) imitate a legitimate service. Another caveat is that Firefox still supports the WebRequestBlocking permission for V3 extensions. Testing in Waterfox obtained the same results.
III-B Extension API Permissions’ Landscape
@@
The above examples highlight how certain extension APIs can be leveraged to steal and modify cookies (including session cookies). To determine the prevalence of these permissions we crawled Chrome’s and Firefox’s extension repositories: the Chrome Web Store [70] and addons.mozilla.org (AMO) [71]. As of March 2024, we identified 122,951 Chrome and 37,783 Firefox extensions. However, of the extensions identified, we could only pull the manifests of 119,152 Chrome and 37,631 Firefox extensions for analysis. Appendix A discusses how we scraped each extension in greater detail. We also identified several Chrome Apps and Android extensions. As Chrome Apps are deprecated on all systems except ChromeOS (where the deprecation is ongoing) [72] and only 22 extensions were marked as Andriod, we exclude them from our main analysis. Instead, we analyze the discovered Apps and Android variants in Appendix B. All extensions and manifests analyzed are available in our public repository [43].
@@
We parsed each extension’s manifest for three attributes. First, we obtained the manifest version from the manifest_version tag. Next, we checked the permissions entry to find if the extension used any of the three APIs discussed in Section III-A. Finally, we searched the manifest for the strings “$<$all_urls$>$”, “*://*/*”, “https://*/*”, and “http://*/*”. These strings allow an extension to access every page (first two strings), all HTTPS pages, and all HTTP pages in the browser, respectively [73]. While the host_permissions key lists the sites an extension interacts with, the entry is optional [74, 75]. DeclarativeNetRequest and content scripts can specify hosts within their own manifest entries. Thus an extension without a host_permissions entry may still have access to cookies. For this reason, we search the whole manifest for the above strings rather than just the host_permissions entry. Detecting these host permissions in an extension indicates that at least a portion of it has this access. Similarly, not all extensions with “$<$all_urls$>$” behavior use this tag. As such, we treat “$<$all_urls$>$”, “*://*/*”, and the combination of “https://*/*” and “http://*/*” as equivalent. Table I summarizes our results. We also calculate the number of users affected by each API as reported by each extension store (depicted in Table II).
@@
We found that 59.9% of extensions are written using manifest V2. However, V2 extensions are much more prevalent in Firefox accounting for 92.8% of all Firefox extensions. Chrome extensions are more evenly split with 49.5% still using the V2 format. This disparity is likely due to Chrome’s (at the time of testing proposed and now deployed) deprecation of the V2 format [63] whereas Firefox will continue to support both versions [64].
@@
For host permissions, we found that 35.2% of extensions analyzed define broad host permissions. Unlike manifest versioning, this ratio is consistent across Firefox (35.6%) and Chrome (35.3%). Nearly all of these extensions have the “$<$all_urls$>$” behavior. This accounts for 34.2%, 35%, and 33.9% of all, Firefox, and Chrome extensions, respectively. 0.2% of extensions declared permission for all HTTP sites only and 1% declared permission for all HTTPS sites alone.
@@
7,882 extensions (5% of extensions analyzed) used the Cookies API, reaching a combined user base of 422,847,233 users. Note that this combined user base represents an upper bound, as a single user may have multiple extensions and thus be counted multiple times. Similarly, it is unclear whether extension repositories measure users by installations or actual user accounts. More extensions used the WebRequest API. Specifically, 11,839 (7.6%) were able to monitor network traffic and 6,978 extensions (4.5%) also declared the WebRequestBlocking permission and could edit HTTP headers. Again while limited extensions declare these permissions, they have nearly one billion users (994,703,506) with 533,244,862 users simultaneously affected by the WebRequestBlocking permission. The DeclarativeNetRequest API is the least used, only appearing in 2,114 extensions (1.3%). Similarly, it is almost non-existent in Firefox appearing in only 66 extensions (0.17% of all Firefox extensions). This low adoption is likely due to the transition from V2 to V3 formatting (at least in Chrome), as the DeclarativeNetRequest API is only available to V3 extensions. In that regard, it appeared in 3.4% of V3 extensions found and has a combined 353,734,697 users.
@@
Each API is limited by the extension’s host permissions. Hence, we also examined how many of the above extensions also have “$<$all_urls$>$” host permissions. Table III summarizes these findings. 54.8% of extensions with Cookies API permissions have access to all cookies in the browser. Similarly, 65.2% of WebRequest extensions and 63.7% of WebRequestBlocking extensions can read and modify any requests in the browser. 67.9% of manifests with DeclarativeNetRequest permissions can also modify any network requests. Further, 21.3% of these extensions (14.5% all declarative extensions) also contain WebRequest permissions, allowing them to view any request (despite contradicting the privacy-preserving philosophy of the declarative API).
@@
Our analysis shows that widespread extensions can access users’ cookies. Similarly, these extensions have a combined user base in the order of hundreds of millions. Thus if compromised (or malicious), these extensions pose a significant threat to cookies and session security.
IV CREAM Overview
@@
Motivated by the results presented in Section III, we advocate for browser controls to allow selective prevention or detection (when prevention is not possible due to application requirements) of cookie theft and illegal cookie modifications made by extensions. To realize these goals we propose CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modification to the browser’s trust model. This Section focuses on CREAM’s concepts and protocol. We discuss browser support for CREAM in Section V.
@@ H $=HMAC(Report_{C_{i}},K_{browser})$
IV-A Threat Model
We extend the active network attacker model described in prior work [18, 19, 22]. The model assumes an Adversary ($\sf{\mathcal{A}dv}$) can intercept, modify, and inject HTTP traffic on the network. It also assumes $\sf{\mathcal{A}dv}$ controls a malicious website. In addition, we assume that $\sf{\mathcal{A}dv}$ controls a malicious extension in the user’s browser with arbitrary API permissions and can run on all pages in the browser (e.g. has the “$<$all_urls$>$” host permission). As such, $\sf{\mathcal{A}dv}$ can also inject content scripts (malicious JS) into all pages. $\sf{\mathcal{A}dv}$’s goal is to modify or steal a user’s cookies through attacks such as cookie hijacking and cookie fixation and gain access to the user’s session.
IV-B Rationale
CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly is a 1-byte flag indicating whether a cookie should be visible to extensions. If set, the cookie is only accessible to the browser. Therefore neither JS nor extensions can access these cookies. BrowserOnly also prevents accesses by JS to ensure an extension cannot access cookies through a content script. The BrowserOnly attribute only protects cookies within the browser. To prevent BrowserOnly cookies from being stolen on the network they should also be marked Secure, which is standard.
For cases where cookies must be accessible for websites and extensions to function properly (where BrowserOnly can not be used), we define the Monitored attribute. The Monitored attribute holds a server-defined message that is stored by and only accessible to the browser. The browser also logs all changes made to Monitored cookies. Later, when sending a Monitored cookie to the server, the browser attaches an authenticated report to the request that contains the server-defined message, the cookie’s current settings, and any logged changes. With this, the server can validate the changes made to the cookie and detect if it is stolen.
Since Monitored cookies must be accessible, we cannot prevent them from being stolen. Instead, we make stolen cookies unusable (e.g., for authentication). Figure 2 depicts the protocol to create (part (A)) and use (part (B)) Monitored cookies. The browser generates a symmetric key for each site it visits (step A.1 in Figure 2) and attaches it to all outbound HTTPS requests for that site (step A.2). Note that in this context, a site refers to the protocol and registry controlled domain name of a resource [20, 76] (e.g., https://example.com). The generated key allows a site to identify the browser in subsequent communication. However, by creating a new key per site the keys cannot be used to track users across sites. When creating a Monitored cookie, the server builds the Monitored attribute’s message. This message must include the cookie’s name, value, and the received browser key. The cookie’s name-value pair ties the attribute to the specific cookie while the browser key ties the whole cookie to a single browser. The server can also include any other data necessary for later cookie validation. The server then uses a secret to symmetrically encrypt the message and uses the resulting ciphertext as the Monitored attribute’s value (step A.3). This authenticated encryption ensures the attribute was generated by the server, is only readable to the server, and cannot be altered.
@@ IV-B Rationale
When sending the Monitored cookie back to the server, the browser builds its corresponding Monitored report, including an HMAC produced with the browser’s key for that site, and attaches it to the request (steps B.1-B.3). Upon receipt, the server decrypts the Monitored message with its secret key (B.4), validates that it is the correct message for the received cookie (B.5), and verifies the HMAC using the key stored in the Monitored attribute (B.6). If the cookie was stolen extension, the HMAC verification would fail as $\sf{\mathcal{A}dv}$ would not know the browser’s key (as long as the browser is itself trusted – recall Section IV-A). Similarly, attempts to change the server-defined Monitored message would cause the decryption/validation to fail. As a result, Monitored cookies can only be used by the legitimate browser.
As detection relies on the symmetric key sent to the server in the initial request, the key can only be sent over secure (HTTPS) channels. This ensures the key is always encrypted and inaccessible to $\sf{\mathcal{A}dv}$ over the network. Otherwise, if not encrypted $\sf{\mathcal{A}dv}$ could steal the key, forge authenticated Monitored reports, and use stolen Monitored cookies. However, Monitored cookies themselves can be sent over any channel as the Monitored attribute is always encrypted.
V CREAM Architectural Design
As discussed above, the BrowserOnly and Monitored attributes require architectural support from the browser to restrict/monitor extensions’ actions. Aside from creating the new cookie attributes, this also requires support from the Cookie Monster and Network Service. As each attribute functions independently, we discuss their required support separately.
V-A BrowserOnly Attribute
As discussed in Section IV, BrowserOnly is a 1-byte flag that prevents a cookie from being accessed by JS and extensions. Thus the Cookie Monster must be updated to filter on this attribute. Recall from Section II-C2 that there are two ways to fetch cookies from the Cookie Monster, GetAll and GetList, and that both GetList and Set filter cookies based on a supplied set of options. As such, we create a new option to exclude BrowserOnly cookies from GetList and Set calls. This option prevents BrowserOnly cookies from being returned or overwritten in most cases. However, as GetAll never filters its results, BrowserOnly cookies are still returned. To address this, we create a new endpoint (GetAllExt) that returns all non-BrowserOnly cookies in the Cookie Monster. These changes are then propagated to the Cookie Manager API and its respective implementations. Finally, the Cookies extension API is updated to use GetAllExt instead of GetAll and the option to exclude BrowserOnly is added to the necessary accesses. This blocks all extension and JS accesses to BrowserOnly cookies through the Cookie Monster. Note that cookie deletion in response to storage restrictions (“purges”) can be conditioned on cookies’ priorities and security attributes [62]. Therefore, we also configure the Cookie Monster to not evict BrowserOnly cookies in these scenarios, thus mitigating eviction attacks [5, 77].